# Congruent modular forms and anticyclotomic Iwasawa theory

### v1

@@ 4.1 $t$-expansion of $p$-adic modular forms
. Denote by $\widehat{\mathbb{Q}_{p}^{nr}}$ the completion of the maximal unramified extension of $\mathbb{Q}_{p}$ and let $\mathcal{W}$ be its ring of integers. Let $\text{Ig}(N)_{/\mathcal{W}}$ be the Igusa scheme over $\mathcal{W}$, and let $\mathbf{x}=[(A_{0},\eta)]\in\text{Ig}(N)(\bar{\mathbb{F}}_{p})$ where $A_{0}$ is an elliptic curve over ${\bar{\mathbb{F}}_{p}}$ and $\eta:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow A_{0}[N]\oplus A_{0}[p^{\infty}]$ is a $\Gamma_{1}(Np^{\infty})$-level structure. Let $\hat{S_{\mathbf{x}}}$ be the local deformation space of $\mathbf{x}$, which represents the functor
$R\longmapsto\{\text{deformations of $A_{0}$ to $R$}\}$
@@
We denote by $q_{\mathcal{A}}$ the pairing corresponding to the isomorphism class $[\mathcal{A}_{/R}]$. As remarked in [5, Section 3.1], $\eta_{p}$ determines a point $P_{\mathbf{x}}\in T_{p}(A_{0}^{t})$ via the Weil pairing, which gives the canonical Serre-Tate coordinates $t:\hat{S_{\mathbf{x}}}\rightarrow\hat{\mathbf{G}}_{m}$ as
@@
[…]
@@
Following the notation of [5, p. 8], for a continuous function $\phi:\mathbb{Z}_{p}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}$, we define $(f\otimes\phi)(t)\in\mathcal{O}_{\mathbb{C}_{p}}\llbracket t-1\rrbracket$ by
[…]
For a classical newform $f$ of weight $2r$ in $S_{2r}^{\text{new}}(\Gamma_{0}(N))$, its Fourier coefficients $\{a_{n}(f)\}_{n>0}$ generate a number field $L$. We may enlarge $\mathcal{W}$ to be the ring of integers of the compositum $\widehat{\mathbb{Q}^{nr}_{p}}\cdot L_{p}$, so that both $\widehat{f}$ and $\widehat{f^{\flat}}$ are $p$-adic modular forms over $\mathcal{W}$. Note that $\mathcal{W}$ is still a complete discrete valuation ring with residue field $\bar{\mathbb{F}}_{p}$. We may then define the $t$-expansions $\widehat{f}(t),\widehat{f^{\flat}}(t)\in\mathcal{W}\llbracket t-1\rrbracket$ as above.
@@ 4.3 Anticyclotomic $p$-adic $L$-function
For a positive integer $c=c_{0}p^{n}$ where $\gcd(c_{0},p)=1$, let $\mathfrak{a}$ be a fractional ideal of $\mathcal{O}_{c}=\mathbb{Z}+c\mathcal{O}_{K}$ and $[(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\in\text{Ig}(N)(K[c])$ be the corresponding CM point on the Igusa scheme discussed in Section 3. Let $\widetilde{\Gamma}_{K}:=\operatorname{Gal}(K[p^{\infty}]/K)$ be the Galois group of the compositum of ring class fields of $K$ with $p^{th}$-power conductor over $K$.
Following [5, p.12], let $\mathfrak{a}\subset\mathcal{O}_{c_{0}}$ be a fractional ideal prime with $Np$ and let $t_{\mathfrak{a}}$ be the canonical Serre-Tate coordinate of $\widehat{f^{\flat}}$ around the reduction $\mathbf{x}_{\mathfrak{a}}=[(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\otimes_{\mathcal{W}}\bar{\mathbb{F}}_{p}$ of $[(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\in\text{Ig}(N)(K[c_{0}])$. Finally, set
@@
where $N(\mathfrak{a})=c^{-1}\#(\mathcal{O}_{c_{0}}/\mathfrak{a})$ ([5, Section 3.2]).
@@
[5, Definition 3.7] Let $c_{0}\geq 1$ be a positive integer such that $(c_{0},pN)=1$ and let $\psi$ be an anticyclotomic Hecke character of infinity type $(r,-r)$ of conductor $c_{0}\mathcal{O}_{K}$. Define $\mathscr{L}_{\mathfrak{p},\psi}(f)$ on $\widetilde{\Gamma}$ to be the $p$-adic measure on $\widetilde{\Gamma}$ given by
@@
for every continuous function $\rho:\widetilde{\Gamma}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}.$ We can also view $\mathscr{L}_{\mathfrak{p},\psi}(f)$ as an element in the semi-local ring $\mathcal{W}\llbracket\widetilde{\Gamma}\rrbracket$. It is known that $\mathscr{L}_{\mathfrak{p},\psi}(f)\neq 0$ [5, Theorem 3.9].
@@ 5 Congruent modular forms
As before, we denote by $f_{1}\in S_{2r_{1}}^{\text{new}}(\Gamma_{0}(N_{1})),f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ normalized Hecke eigenforms of weight $2r_{1}$, $2r_{2}$ and levels $N_{1},N_{2}$, respectively. Suppose that both $f_{1}$ and $f_{2}$ satisfy Hypothesis (1). Then there exist ideals $\mathfrak{N}_{1}$, $\mathfrak{N}_{2}$ in $\mathcal{O}_{K}$ such that $N_{1}\mathcal{O}_{K}=\mathfrak{N}_{1}\overline{\mathfrak{N}_{1}}$ and $N_{2}\mathcal{O}_{K}=\mathfrak{N}_{2}\overline{\mathfrak{N}_{2}}$. Further assume that for every $\ell\mid\gcd(N_{1},N_{2})$, one has $\gcd(\ell,\mathfrak{N}_{1})=\gcd(\ell,\mathfrak{N}_{2})$ so that $\mathcal{O}_{K}/\operatorname{lcm}(\mathfrak{N}_{1},\mathfrak{N}_{2})\simeq\mathbb{Z}/\operatorname{lcm}(N_{1},N_{2})\mathbb{Z}$.
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}}$, $f_{2}\in S_{2r_{2}}^{\text{new}}(\Gamma_{0}(N_{2}))$ have the same level $N_{1}=N_{2}$. Let $L$ be a number field containing ($\mathbb{Q}(\{a_{n}(f_{1}),a_{n}(f_{2})\}_{n>0})$ and let $\varpi$ be a uniformizer of $\mathcal{O}_{L_{p}}$. Suppose that $a_{n}(f_{1})\equiv a_{n}(f_{2})\pmod{\varpi^{m}\mathcal{O}_{L_{p}}}$ for every $n$. Then we have the congruences $f_{1}^{\flat}\equiv f_{2}^{\flat}\pmod{\varpi^{m}\mathcal{O}_{L_{p}}}$, $\widehat{f_{1}^{\flat}}\equiv\widehat{f_{2}^{\flat}}\pmod{\varpi^{m}\mathcal{O}_{L_{p}}}$ between $p$-adic modular forms, and $\mathscr{L}_{\mathfrak{p}}(f_{1})\equiv\mathscr{L}_{\mathfrak{p}}(f_{2})\pmod{\varpi^{m}\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket}.$
@@
The congruences between $p$-adic modular forms follow from the $q$-expansion principle [12, Corollary 3.5]. We show that $\mathscr{L}_{\mathfrak{p}}(f_{1})(\rho)\equiv\mathscr{L}_{\mathfrak{p}}(f_{2})(\rho)\pmod{\varpi^{m}\mathcal{O}_{\mathbb{C}_{p}}}$ for every continuous map $\rho:\widetilde{\Gamma}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}^{\times}$, and the congruence $\mathscr{L}_{\mathfrak{p}}(f_{1})\equiv\mathscr{L}_{\mathfrak{p}}(f_{2})\pmod{\varpi^{m}\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket}$ follows by the same argument as [20, Theorem (1.10)]. Let $\chi_{\text{cyc}}:G_{\mathbb{Q}}\rightarrow\mathbb{Z}_{p}^{\times}\subset\mathcal{O}_{L_{p}}^{\times}$ be the cyclotomic character and let $\rho_{f_{i}}$ be the Weil-Deligne representation attached to $f_{i}$ for $i\in\{1,2\}$. Since $\det(\rho_{f_{i}})=\chi_{\text{cyc}}^{2r_{i}-1}$, we have the congruence
$\chi_{\text{cyc}}^{2r_{1}-1}\equiv\chi_{\text{cyc}}^{2r_{2}-1}\pmod{1+\varpi^{m}\mathcal{O}_{L_{p}}}.$
Suppose that $\varpi^{m}\mathcal{O}_{L_{p}}\cap\mathbb{Z}_{p}=p^{m^{\prime}}\mathbb{Z}_{p}$, then the congruence above actually holds in $(\mathbb{Z}_{p}/p^{m^{\prime}}\mathbb{Z}_{p})^{\times}\subset(\mathcal{O}_{L_{p}}/\varpi^{m}\mathcal{O}_{L_{p}})^{\times}:$
@@
We define $\psi$ to be the Hecke character such that $\psi(\mathfrak{a})=\left(\frac{N(\mathfrak{a})}{p}\right)$ for prime-to-$p$ fractional ideals $\mathfrak{a}$ of $K$. Then $\psi$ is an anticyclotomic Hecke character of order $2$ and conductor $p$, and $\psi_{\mathfrak{p}}:\mathcal{O}_{K,\mathfrak{p}}^{\times}\rightarrow\{\pm 1\}$ is the Legendre symbol $\left(\frac{\cdot}{p}\right)$. We may now rewrite the congruence (5.1) as
@@ 5.1 Hecke operators at $p$ in Serre-Tate coordinates
We recall some Hecke operators in terms of the complex uniformization of Igusa schemes. Let $\mathfrak{a}$ be a fractional ideal of $\mathcal{O}_{K}$ and let $x_{\mathfrak{a}}=[(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]=[\vartheta,\rho(\overline{a})^{-1}\zeta_{c}]$ (see Section 3). For $z\in\mathbb{Q}_{p}$, we define $\mathfrak{n}(z):=\begin{pmatrix}1&z\\ 0&1\end{pmatrix}\in\text{GL}_{2}(\mathbb{Q}_{p})\subset\text{GL}_{2}(\widehat{\mathbb{Q}})$ and let $x_{\mathfrak{a}}\ast\mathfrak{n}(z):=[\vartheta,\rho(\overline{a})^{-1}\zeta_{c}\mathfrak{n}(z)]$ under the action of $\text{GL}_{2}(\widehat{\mathbb{Q}})$ on $\text{Ig}(N)(\mathbb{C})$.
By [5, Proposition 3.3], for a primitive Dirichlet character $\phi:(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}^{\times}$, the integral in Definition 4.1 can be written as
@@
In [4, Proposition 6.4], the author discusses the moduli interpretations of $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})$ for $u\in(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$ as quotients of $A_{\mathfrak{a}}$ by certain rank-$p^{n}$ subgroup schemes of $A_{\mathfrak{a}}[p^{\infty}]$. Moreover we have $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})\otimes\bar{\mathbb{F}}_{p}=x_{\mathfrak{a}}\otimes\bar{\mathbb{F}}_{p}$, and the Serre-Tate coordinate of $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})$ is given by $t_{\mathfrak{a}}(x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n}))=\zeta_{p^{n}}^{-u\mathbb{N}(\mathfrak{a})^{-1}\sqrt{-D_{K}}^{-1}}$ [5, Lemma 3.2].

### v2

@@ 4.1 $t$-expansion of $p$-adic modular forms
Recall that $\widehat{\mathbb{Q}_{p}^{nr}}$ is the completion of the maximal unramified extension of $\mathbb{Q}_{p}$ and $\mathcal{W}$ is its ring of integers. Let $\text{Ig}(N)_{/\mathcal{W}}$ be the Igusa scheme over $\mathcal{W}$, and let $\mathbf{x}=[(A_{0},\eta)]\in\text{Ig}(N)(\bar{\mathbb{F}}_{p})$ where $A_{0}$ is an elliptic curve over $\bar{\mathbb{F}}_{p}$} and $\eta:\mu_{N}\oplus\mu_{p^{\infty}}\hookrightarrow A_{0}[N]\oplus A_{0}[p^{\infty}]$ is a $\Gamma_{1}(Np^{\infty})$-level structure. Let $\hat{S_{\mathbf{x}}}$ be the local deformation space of $\mathbf{x}$, which represents the functor
[…]
@@
We denote by $q_{\mathcal{A}}$ the pairing corresponding to the isomorphism class $\mathcal{A}_{/R}]$. As remarked in [3, Section 3.1], $\eta_{p}$ determines a point $P_{\mathbf{x}}\in T_{p}(A_{0}^{t})$ via the Weil pairing, which gives the canonical Serre-Tate coordinates $t:\hat{S_{\mathbf{x}}}\rightarrow\hat{\mathbf{G}}_{m}$ as
@@
$\varphi\mapsto\Phi_{\varphi}(t)=\sum_{n=0}^{\infty}\left(\int_{\mathbb{Z}_{p}}{x\choose n}d\varphi(x)\right)(t-1)^{n}=\int_{\mathbb{Z}_{p}}t^{x}d\varphi(x),$
@@
Following the notation of [3, p. 8], for a continuous function $\phi:\mathbb{Z}_{p}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}$, we define $(f\otimes\phi)(t)\in\mathcal{O}_{\mathbb{C}_{p}}\llbracket t-1\rrbracket$ by
$(f\otimes\phi)(t)=\int_{\mathbb{Z}_{p}}\phi(x)t^{x}df=\sum_{n\geq 0}\int_{\mathbb{Z}_{p}}\phi(x){x\choose n}df(x)\cdot(t-1)^{n}.$
For a classical newform $f$ of weight $2r$ in $S_{2r}^{\text{new}}(\Gamma_{0}(N))$, its Fourier coefficients $\{a_{n}(f)\}_{n>0}$ lie in a $p$-adic field $F$. We may enlarge $\mathcal{W}$ to be the ring of integers of the compositum $\widehat{\mathbb{Q}^{nr}_{p}}\cdot F$, so that both $\widehat{f}$ and $\widehat{f^{\flat}}$ are $p$-adic modular forms over $\mathcal{W}$. Note that $\mathcal{W}$ is still a complete discrete valuation ring with residue field $\bar{\mathbb{F}}_{p}$ and define the $t$-expansions $\widehat{f}(t),\widehat{f^{\flat}}(t)\in\mathcal{W}\llbracket t-1\rrbracket$ as above.
@@ 4.3 Anticyclotomic $p$-adic $L$-function
For a positive integer $c=c_{0}p^{n}$ where $\gcd(c_{0},p)=1$, let $\mathfrak{a}$ be a fractional ideal of $\mathcal{O}_{c}=\mathbb{Z}+c\mathcal{O}_{K}$ and $(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\in\text{Ig}(N)(K[c])$ be the corresponding CM point on the Igusa scheme discussed in Section 3. Let $\widetilde{\Gamma}_{K}:=\operatorname{Gal}(K[p^{\infty}]/K)$ be the Galois group of the compositum of ring class fields of $K$ with $p^{th}$-power conductor over $K$.
Following [3, p.12], let $\mathfrak{a}\subset\mathcal{O}_{c_{0}}$ be a fractional ideal prime with $Np$ and let $t_{\mathfrak{a}}$ be the canonical Serre-Tate coordinate of $\widehat{f^{\flat}}$ around the reduction $\mathbf{x}_{\mathfrak{a}}=[(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\otimes_{\mathcal{W}}\bar{\mathbb{F}}_{p}$ of $(A_{\mathfrak{a}},\eta_{\mathfrak{a}})]\in\text{Ig}(N)(K[c_{0}])$. Finally, set
@@
where $N(\mathfrak{a})=c^{-1}\#(\mathcal{O}_{c_{0}}/\mathfrak{a})$ ([3, Section 3.2]).
@@
[3, Definition 3.7] Let $c_{0}\geq 1$ be a positive integer such that $(c_{0},pN)=1$ and let $\psi$ be an anticyclotomic Hecke character of infinity type (r,r)(r,-r) of conductor $c_{0}\mathcal{O}_{K}$. Define $\mathscr{L}_{\mathfrak{p},\psi}(f)$ on $\widetilde{\Gamma}$ to be the $p$-adic measure on $\widetilde{\Gamma}$ given by
@@
for every continuous function $\rho:\widetilde{\Gamma}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}.$ We can also view $\mathscr{L}_{\mathfrak{p},\psi}(f)$ as an element in the semi-local ring $\mathcal{W}\llbracket\widetilde{\Gamma}\rrbracket$. It is known that $\mathscr{L}_{\mathfrak{p},\psi}(f)\neq 0$ [3, Theorem 3.9].
@@ 5 Congruent modular forms
Let $f_{1}\in S_{2r_{1}}^{\text{new}}(\Gamma_{0}(N_{1})),f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ be normalized Hecke eigenforms of weight $2r_{1}$, $2r_{2}$ and levels $N_{1},N_{2}$, respectively. Suppose that both $f_{1}$ and $f_{2}$ satisfy Hypothesis (1). Then there exist ideals $\mathfrak{N}_{1}$, $\mathfrak{N}_{2}$ in $\mathcal{O}_{K}$ such that $N_{1}\mathcal{O}_{K}=\mathfrak{N}_{1}\overline{\mathfrak{N}_{1}}$ and $N_{2}\mathcal{O}_{K}=\mathfrak{N}_{2}\overline{\mathfrak{N}_{2}}$. Further assume that for every $\ell\mid\gcd(N_{1},N_{2})$, one has $\gcd(\ell,\mathfrak{N}_{1})=\gcd(\ell,\mathfrak{N}_{2})$ so that $\mathcal{O}_{K}/\operatorname{lcm}(\mathfrak{N}_{1},\mathfrak{N}_{2})\simeq\mathbb{Z}/\operatorname{lcm}(N_{1},N_{2})\mathbb{Z}$.
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}}$, $f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ have the same level $N_{1}=N_{2}$. Let $F$ be a $p$-adic field containing $\mathbb{Q}(\{a_{n}(f_{1}),a_{n}(f_{2})\}_{n>0})$ and let $\varpi$ be a uniformizer of $\mathcal{O}_{F}$. Suppose that $a_{n}(f_{1})\equiv a_{n}(f_{2})\pmod{\varpi^{m}\mathcal{O}_{F}}$ for every $n$. Then there are congruences $f_{1}^{\flat}\equiv f_{2}^{\flat}\pmod{\varpi^{m}\mathcal{O}_{F}}$, $\widehat{f_{1}^{\flat}}\equiv\widehat{f_{2}^{\flat}}\pmod{\varpi^{m}\mathcal{O}_{F}}$ between $p$-adic modular forms, and
@@
The congruences between $p$-adic modular forms follow from the $q$-expansion principle [12, Corollary 3.5]. We show that $\mathscr{L}_{\mathfrak{p}}(f_{1})(\rho)\equiv\mathscr{L}_{\mathfrak{p}}(f_{2})(\rho)\pmod{\varpi^{m}\mathcal{O}_{\mathbb{C}_{p}}}$ for every continuous map $\rho:\widetilde{\Gamma}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}^{\times}$, and the congruence $\mathscr{L}_{\mathfrak{p}}(f_{1})\equiv\mathscr{L}_{\mathfrak{p}}(f_{2})\pmod{\varpi^{m}\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket}$ follows by the same argument as [20, Theorem (1.10)]. Let $\chi_{\text{cyc}}:G_{\mathbb{Q}}\rightarrow\mathbb{Z}_{p}^{\times}\subset\mathcal{O}_{F}^{\times}$ be the cyclotomic character and let $\rho_{f_{i}}$ be the Weil-Deligne representation attached to $f_{i}$ for $i\in\{1,2\}$. Since $\det(\rho_{f_{i}})=\chi_{\text{cyc}}^{2r_{i}-1}$, we have the congruence
$\chi_{\text{cyc}}^{2r_{1}-1}\equiv\chi_{\text{cyc}}^{2r_{2}-1}\pmod{1+\varpi^{m}\mathcal{O}_{F}}$.
Suppose that $\varpi^{m}\mathcal{O}_{F}\cap\mathbb{Z}_{p}=p^{m^{\prime}}\mathbb{Z}_{p}$, then the congruence above actually holds in $(\mathbb{Z}_{p}/p^{m^{\prime}}\mathbb{Z}_{p})^{\times}\subset(\mathcal{O}_{F}/\varpi^{m}\mathcal{O}_{F})^{\times}:$
@@
Define $\psi$ as the Hecke character such that $\psi(\mathfrak{a})=\left(\frac{N(\mathfrak{a})}{p}\right)$ for prime-to-$p$ fractional ideals $\mathfrak{a}$ of $K$. Then $\psi$ is an anticyclotomic Hecke character of order $2$ and conductor $p$, and $\psi_{\mathfrak{p}}:\mathcal{O}_{K,\mathfrak{p}}^{\times}\rightarrow\{\pm 1\}$ is the Legendre symbol $\left(\frac{\cdot}{p}\right)$. We may now rewrite the congruence (5.1) as
@@ 5.1 Hecke operators at $p$ in Serre-Tate coordinates
[…]
By [3, Proposition 3.3], for a primitive Dirichlet character $\phi:(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}\rightarrow\mathcal{O}_{\mathbb{C}_{p}}^{\times}$, the integral in Definition 4.1 can be written as
@@
In [5, Proposition 6.4], the author discusses the moduli interpretations of $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})$ for $u\in(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$ as quotients of $A_{\mathfrak{a}}$ by certain rank-$p^{n}$ subgroup schemes of $A_{\mathfrak{a}}[p^{\infty}]$. Moreover we have $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})\otimes\bar{\mathbb{F}}_{p}=x_{\mathfrak{a}}\otimes\bar{\mathbb{F}}_{p}$, and the Serre-Tate coordinate of $x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n})$ is given by
$t_{\mathfrak{a}}(x_{\mathfrak{a}}\ast\mathfrak{n}(up^{-n}))=\zeta_{p^{n}}^{-u\mathbb{N}(\mathfrak{a})^{-1}\sqrt{-D_{K}}^{-1}}$
according to [3, Lemma 3.2].